# Augmented Lagrangian Solvers for Poroelasticity with Fracture Contact Mechanics

### v1

@@ 4.2 Three-dimensional simulations of fluid injection
Next, the solvers will be tested on a three-dimensional problem. As in the previous section, we aim to simulate hydraulic stimulation of critically stressed fractures by an injection well. The domain is the box $(0\text{m},2000\text{m})\times(0\text{m},2000\text{m})\times(-3000\text{m},-1000\text{m})$, and contains a network of three intersecting, elliptic-shaped fractures, as shown in figure 7. For the hydraulic boundary conditions, we impose hydrostatic pressure conditions on all external boundaries, given by $p=\rho^{f}_{\mathrm{ref}}gz$, with $z$ being the depth and $g$ the gravitational acceleration. For the mechanical boundary conditions, we fix the bottom boundary (zero displacement), while for the remaining boundaries, we impose compressive tractions, with a similar anisotropy as used in section 4.1. Namely, the corresponding non-zero stress tensor values are:
@@
The rest of the simulation setup is similar to that of section 4.1. We will keep the centermost cell of one of the fractures at a constant higher pressure to mimic an injection well (indicated by the vertical black line in figure 7). We test three different injection pressures, by adding 1, 8 and 10 MPa to the background hydrostatic pressure in the injection cell, respectively. The dilation angle is set to $5^{\circ}$ for all simulations. Initial values of the variables are found by imposing a constant hydrostatic pressure field and solving only the purely mechanical subproblem. Again we take a single time step, of size 0.1 days.
Figure 7: Left: Fracture network geometry used in the three-dimensional simulations, and the position of the injection well, marked by the vertical black line. Also shown are the two-dimensional grid cells, as well as parts of the three-dimensional grid. Finally, the contact states (as defined by (28)) after the initialization phase are also partially shown, where fracture slip is measured relative to a configuration with no deformation. A total of 702 cells are sticking, and 112 cells are slipping. Right: Contact states after a time step of 0.1 days, for the injection pressure of 30 MPa. A total of $5$ cells are open, 179 cells are sticking, and 580 cells are slipping. As before, fracture slip is measured relative to the state after the initialization
The performance of the solvers (but now excluding IRM) for different values of the augmentation parameter and for the three different injection pressures, is shown in figure 8. We observe relatively similar behavior to the two-dimensional case. Increasing the injection pressure causes an increase in the iteration counts, as well as the number of cases not converging. For the two lowest injection pressures, we observe GNM being quite consistent for the lower $c$-values, but failing to converge for the higher values, while GNM-RM converges across a larger range of $c$-values, including the higher ones. These are the same results as observed for the lower injection pressures in the two-dimensional example (figures 3(a)-3(c)).
However, we once again see the pattern being broken for the most difficult case, with the convergence behavior being more erratic. In particular, GNM-RM fails to converge in all cases. Investigation of the residual norms for a particular case (we checked for $c=10^{3}$) reveals that Newton once again enters a cycle. This cycle is avoided by GNM-RM, but its residual norm nonetheless fails to fall below the threshold within 100 iterations. We conclude from these observations that, while GNM-RM displays in many cases added robustness with respect to the augmentation parameter, this is not necessarily always the case.
Figure 8: Iteration counts for one time step for the example of section 4.2 with different injection pressures. NC and Div denote simulations where the solver did not converge within 100 iterations or diverged to infinity, respectively
@@ 5 Conclusion
We have investigated the performance of three different nonlinear solvers, all based on the augmented Lagrangian formulation of frictional contact mechanics, for solving advanced problems of coupled flow and deformation in fractured porous media, including fracture contact mechanics. This included two classical solvers, namely the generalized Newton method (GNM) using complementarity functions, and the implicit return map method (IRM). In addition, we have proposed a new nonlinear solver, which essentially combined the two classical solvers. Specifically, the new nonlinear solver incorporated the return map into the generalized Newton method as a postprocessing step after every iteration. It was subsequently named the ”generalized Newton method with a return map” (GNM-RM). The purpose of the return map in GNM-RM was to stabilize the generalized Newton method, which is in contrast to classical return map methods like IRM, where its purpose is to ensure the feasibility of the contact tractions.
The performance of the three solvers was assessed through a series of numerical experiments, both two- and three-dimensional, which were designed to mimic hydraulic stimulation of a fractured geothermal reservoir. Key questions included the assessment of the robustness and sensitivity of the performance with respect to the augmentation parameter, as well as grid resolution. The experiments indicated that convergence issues are easily encountered, even for relatively simple fracture geometries. For the simplest numerical experiments, the two classical solvers, GNM and IRM, behaved more or less as expected and in accordance with established heuristics from the contact mechanics literature, regarding the augmentation parameter. However, as the experiments became more complex and strongly coupled, the two solvers behaved more erratically, being very sensitive to the value of the augmentation parameter, and to small changes in the input data (as shown by varying the shear dilation angle). In particular, IRM failed to converge in nearly all cases, except for the very simplest experiment. The new solver, GNM-RM, performed in most cases more robustly with respect to the augmentation parameter, and in particular, it handled larger values of the parameter much better than the classical solvers. It was also demonstrated that in some cases where GNM fails to converge due to cycling, the cycle can be avoided by instad using GNM-RM. That being said, convergence issues were also encountered with GNM-RM, particularly in the three-dimensional case, with its behavior again being unpredictable for the most difficult experiments.
In conclusion, for complex problems of mixed-dimensional poromechanics with fracture contact mechanics, we would strongly recommend GNM over IRM. Moreover, GNM-RM can be worthwhile to try if GNM fails to converge, as we have demonstrated that it may converge in cases where GNM fails, and it is easily implemented as a simple modification of GNM. Perhaps more importantly, however, we conclude that convergence issues with nonlinear solvers pose a significant bottleneck for running realistic simulations of fractured geothermal reservoirs, as established solvers will frequently display unpredictable behavior. Clearly, more work is needed on this topic.

### v2

@@ 4.4 Three-dimensional simulations of fluid injection
Next, we investigate the performance of the nonlinear solvers on a set of three-dimensional problems, and investigate to which degree the results from the two-dimensional case carry over to three dimensions. As before, our aim is to simulate hydraulic stimulation of critically stressed fractures by an injection well. The domain is the box $(0\text{m},2000\text{m})\times(0\text{m},2000\text{m})\times(-3000\text{m},-1000\text{m})$, and contains a network of three intersecting, elliptic-shaped fractures, as shown in figure 6. For the hydraulic boundary conditions, we impose hydrostatic pressure conditions on all external boundaries, given by $p=\rho^{f}_{\mathrm{ref}}gz$, with $z$ being the depth and $g$ the gravitational acceleration. For the mechanical boundary conditions, we fix the bottom boundary (zero displacement), while for the remaining boundaries, we impose compressive tractions, with a similar anisotropy as used in section 4.3. Namely, the corresponding non-zero stress tensor values are:
@@
The rest of the simulation setup is similar to that of section 4.3. The pressure in the centermost cell of one of the fractures is momentarily increased at certain points, mimicking an injection well (illustrated by the vertical black line in figure 6). The total simulation time is three hours as in section 4.3, with the pressure profile of the injection cell again following figure 2. The adaptive time-stepping scheme outlined in section 4.2 is employed. The initial state is the equilibrium state with a hydrostatic pressure field. Again we will test the effects of different degrees of nonlinear couplings between flow and mechanics, by using the three different models of table 2.
Figure 6: Fracture network geometry used in the three-dimensional simulations, and the position of the injection well, marked by the vertical black line. Also shown are the two-dimensional grid cells, as well as parts of the three-dimensional grid. The grid contains a total of 16405 cells, including matrix, fracture and interface cells. The average size of the grid cells is seen to be larger near the boundary than on the fractures, indicating grid refinement near the fractures. The different colors on the fractures are used to visually distinguish them, and do not refer to any physical quantity
The resulting contact states and differences in volumetric aperture is shown at various points in the simulations in figure 7. We note that slip around the injection well does not occur until the second phase, when a sufficiently high pressure is reached. Otherwise, we observe qualitatively similar results to the two-dimensional case. The slip events result in an increase of the volumetric aperture (except in the constant-aperture case of model A), due to shear dilation, and the magnitude of the volumetric aperture difference increases from models B to C.
Figure 7: Visualization of the simulations of three-dimensional injection. Results are shown at the end of each simulation phase using model C, and at the end of the final simulation phase using models A and B. Top row: Contact states. Here, slip is measured in a cumulative manner; a cell is regarded as being in a ”slip” state if the fracture is closed and the norm of the tangential displacement jump is positive, using the state before the injection (at $t=0$) as the reference. There is no opening of fractures. Bottom row: Changes in volumetric aperture, relative to the value at $t=0$
Figure 8: Bar plots showing the cumulative number of nonlinear iterations, summed over all time steps, for the example of section 4.4 with the different degrees of nonlinear couplings between flow and mechanics outlined in table 2. The three different colors per bar correspond to the three different phases of the simulation, after each pressure increase in the injection well. Hatched black lines indicate steps where the nonlinear solver failed to converge, and red crosses indicate a simulation that was terminated. The simulation is terminated if one of three cases occur: The number of nonlinear iterations exceed 800, the nonlinear solver failed to converge for the smallest allowed time step, or after 6 unsuccessful recomputation attempts. Finally, the total number of iterations of the linear solver is marked by a black dot
The performance of the nonlinear solvers for different values of the augmentation parameter is shown in figure 8. We observe many of the same patterns as in the two-dimensional simulations. An increase in difficulty occurs from models A through C, as the nonlinear coupling strength is increased. The performance of IRM deteriorates for models B and C, with no cases converging within 800 iterations. In regards to the augmentation parameter, GNM and IRM both fail to converge for the largest three values, while GNM-RM converges in all cases. These results are all very similar to the results of section 4.3.
The behavior of the linear solver also mostly matches that of the two-dimensional simulations. The average number of GMRES iterations per nonlinear iteration slightly decreases with increasing augmentation parameter for IRM, while it is close to constant with respect to the augmentation parameter for GNM and GNM-RM. The same behavior was seen in the two-dimensional simulations of figure 4. A slight difference from the two-dimensional simulations is that for the cases considered here, the return map of GNM-RM does not appear to improve the performance of the linear solver. In nearly all cases of figure 4, GNM-RM used less GMRES iterations per nonlinear iteration than GNM, while in figure 8, this number is about equal for GNM and GNM-RM.
Another notable difference from the two-dimensional simulations is that the fully coupled model is significantly more challenging than the other two models. It is apparent from figure 8(c) that for model C, a considerable spike in difficulty occurs during the third phase, with a large increase in the number of nonlinear iterations. Additionally, GNM fails to converge even for some of the lower $c$-values during this phase (although it is successful after the time step is lowered).
4.5 More challenging three-dimensional simulations
To further explore the challenges of fully coupled three-dimensional simulations and test the limits of GNM and GNM-RM, a final set of simulations has been run using model C. We discard IRM in this final set of simulations, as it has been consistently outclassed by the other two solvers, and instead focus on the comparison between GNM and GNM-RM. The simulation setup is nearly identical to the previous one; the only difference is a reduction in the compressive tractions on the boundary faces aligned with the $y$-direction. The stress field on the boundary is now as follows:
$\sigma_{xx}=-0.6\rho^{s}gz\ \ ,\ \ \sigma_{yy}=-0.4\rho^{s}gz\ \ ,\ \ \sigma_{zz}=-\rho^{s}gz.$ (38)
This reduction in the compressive stress field causes a much larger portion of the fractures to slip, as shown in figure 9, creating more complex nonlinear dynamics that challenge the solvers. Due to the added difficulty of this setup, we will run only the first phase of the injection schedule of figure 2, i.e. we increase the pressure in the injection well by 1MPa at $t=0$, and run the simulation for one hour. The results are reported in figure 10.
Figure 9: Visualization of the simulation of section 4.5, after one hour of simulation time. Left: Fracture contact states. Here, slip is measured in a cumulative manner; a cell is regarded as being in a ”slip” state if the fracture is closed and the norm of the tangential displacement jump is positive, using the state before the injection (at $t=0$) as the reference. There is no opening of fractures. Right: Changes in volumetric aperture, relative to the value at $t=0$
Figure 10: Bar plots showing the cumulative number of nonlinear iterations, summed over all time steps, for the example of section 4.5. Hatched black lines indicate steps where the nonlinear solver failed to converge, and red crosses indicate a simulation that was terminated. The simulation is terminated if one of three cases occur: The number of nonlinear iterations exceed 800, the nonlinear solver failed to converge for the smallest allowed time step, or after 6 unsuccessful recomputation attempts. Finally, the total number of iterations of the linear solver is marked by a black dot.
We observe a significant increase in difficulty for the solvers. The total number of nonlinear iterations is now comparable to the previous three-dimensional simulation with model C (figure 8(c)) for only one third of the simulation time. Moreover, in all cases the nonlinear solvers failed to converge at some point, resulting in a recomputation with a smaller time step. The number of GMRES iterations per nonlinear iteration is also much higher for this example than the previous ones (note the change of scale on the axis for the linear iterations, compared to the previous bar plots). Despite the increased difficulty, GNM-RM remains relatively robust, managing to converge within 800 iterations in all but two cases.
Overall, the results of the three-dimensional simulations of sections 4.4 and 4.5 are relatively similar to the two-dimensional simulations of section 4.3. The increase in difficulty from models A to C is once again clearly seen. IRM is the worst-performing solver also in three dimensions, and deteriorates for models B and C, while GNM-RM is the best-performing one, being the most robust solver with respect to the augmentation parameter. However, we also observe some more unpredictable behavior of the nonlinear solvers for the most difficult case of section 4.5. This is unsurprising, given the high complexity of the fully coupled, three-dimensional model.
@@ 5 Conclusion
We have investigated the performance of three different nonlinear solvers, all based on the augmented Lagrangian formulation of frictional contact mechanics, for solving advanced problems of coupled flow and deformation in fractured porous media, including fracture contact mechanics. This included two well-known solvers from the contact mechanics literature, namely the generalized Newton method (GNM) using complementarity functions, and the implicit return map method (IRM), the latter of which is equivalent to an implicit Uzawa method. In addition, we proposed a third nonlinear solver, which combined features of both GNM and IRM. This solver, named the “generalized Newton method with a return map” (GNM-RM), incorporates the return map into the generalized Newton method as a postprocessing step after every iteration. The purpose of the return map in GNM-RM is to stabilize the generalized Newton method, which is in contrast to classical return map methods like IRM, where its purpose is to ensure the feasibility of the contact tractions.
The performances of the three solvers were assessed through a series of numerical experiments, both two- and three-dimensional, which were designed to mimic hydraulic stimulation of a fractured geothermal reservoir. In the experiments, the degree of nonlinear coupling between flow and mechanics was varied, and different values of the augmentation parameter were tested. The experiments showed that increasing the nonlinear coupling between flow and mechanics significantly adds to the difficulty of the problem, with a higher frequency of non-convergent behavior of the nonlinear solvers, as well as a general increase in the number of nonlinear iterations needed for convergence. GNM and GNM-RM still managed to converge in many cases for the more strongly coupled problems, but IRM performed much worse, barely converging at all except for the cases with very weak couplings. In regards to the augmentation parameter, GNM-RM was the most robust solver, converging across the largest range of values of this parameter. In particular, it handled higher values of the parameter much better than the other two solvers. For the higher values of the augmentation parameter, GNM and IRM would frequently fail to converge, with cycling of the iterations being the central failure more, while GNM-RM would avoid the cycle and converge in nearly all cases. This makes GNM-RM a viable alternative to GNM for the type of problems considered herein.
Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 101002507).